OpenStudy (anonymous):

Using the distance formula to find the distance, to the nearest tenthn between f(9,5) and g(-2,2) show work

OpenStudy (anonymous):

\[d = \sqrt{(x _{2} - x _{1})^{2} + (y _{2} - y _{1})^{2}}\] so d = \[\sqrt{(-2 - 9)^{2} + (2 - 5)^{2}}\] \[d = \sqrt{(-11)^{2} + (-3)^{2}}\] \[d = \sqrt{121 + 9}\] \[d = \sqrt{130}\] so d = use your calculator

OpenStudy (anonymous):

from -2 to 9 is 11 on the x axis from 2 to 5 is 3 on the y axis. use pythagoras and get \[d^2=11^2+3^2=121+9=130\] so \[d=\sqrt{130}\]